A heavy flywheel of mass 15kg and radius 20cm is mounted on an horizontal axle of radius 1cm and negligible mass compared with the flywheel, neglecting friction,find the angular acceleration if a force of 4kg is applied tangentially to the axle
1
Expert's answer
2019-11-07T09:09:24-0500
"I= \frac{t}{a}"

Where I moment of inertion of flywheel, t - torque, a - angular acceleration


"t = FR"

Where R - radius of flywheel, F - force


"I = \frac{M(R-r) ^2}{2}"

Where M- mass of flywheel, r - radius of axle


"a = \frac{2FR }{M(R-r) ^2}= 2.95 (1/s^2)"